OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (anonymous):

one second let me type the equation

OpenStudy (alexandervonhumboldt2):

m offilene for a secom

OpenStudy (anonymous):

\[if f(x)=2x/2x+1 and g(x)=1/2x+1, find f(x)+g(x)\]

OpenStudy (alexandervonhumboldt2):

im back

OpenStudy (alexandervonhumboldt2):

2x/2x+1+1/2x+1 1+1+1+1/2x=1/2x+3

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

next step

OpenStudy (alexandervonhumboldt2):

ok

OpenStudy (anonymous):

next step